OpenStudy (anonymous):

For a fixed voltage, the current A in an electrical circuit is inversely proportional to the reistance R in the circuit, that is, A=k/R, where k is a constant. If the current is 100 amperes when the resistance is 20 ohms, when the resistance is 25 ohms the current is _______ amperes. a. 80 b. 125 c. 120 d. none of these

OpenStudy (amistre64):

plug and play really 100 = k/20 2000 = k A = 2000/25

OpenStudy (amistre64):

a is my vote

OpenStudy (amistre64):

80*25 = ummm....carry the one and cross the i.... 2000